OpenStudy (anonymous):

FIND THE VALUE OF y. 3y+20.....5y-16

OpenStudy (anonymous):

they're together

OpenStudy (anonymous):

the answer is 18 but HOWWW

OpenStudy (anonymous):

ok first equate them, 3y+20=5y-16 then collect y's to one side numbers to the other 3y-5y=-16-20 -2y = -36 then divide by -2 on both sides y=18

OpenStudy (anonymous):

Y DID U SUBTRACT BOTH SIDES.. How did u know to subtract?